A flywheel slows from 600 to 400 rev/min while rotating through 40 revolutions


1) Angular Acceleration of Flywheel

2) The Elapsed During 40 Revolutions


1
Expert's answer
2021-03-18T20:11:33-0400

By definition, the angular acceleration is:



"\varepsilon =\dfrac{\Delta \omega }{ t}"

where "\Delta \omega = 400\space rev/min - 600\space rev/min = -200\space rev/min" is the gain in angular velocity in time "t". On the other hand, the number of revolutions under the constant acceleration are given as follows:


"\varphi = \omega_0t+ \dfrac{\varepsilon t^2}{2}"


where "\omega_0 = 600\space rev/min" is the initial angular velocity. Substituting the expression for "\varepsilon" and expressing "t", obtain:


"t= \dfrac{\varphi}{\omega_0 + \Delta \omega /2}"

since "\varphi = 40", obtain:


"t= \dfrac{40}{600 -200 /2} = 0.08 min"

The acceleration is then:


"\varepsilon = -\dfrac{200\space rev/min}{0.08min} = -2500\space rev/min^2"

Answer. 1) -2500 rev/min, 2) 0.08 min.